# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v2

@@ Abstract.
RAG inference, Acceleration, Prefetching.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: submissionid: 164
1. Introduction
(a) Typical pipeline stages of a RAG application.
(b) Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems. TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU-CPU cooperation.
Figure 1. Overview of RAG pipeline stages and comparison of TeleRAG to baseline systems.
@@ 1. Introduction
Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020). During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and incorporate up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
Modern RAG applications share two key characteristics. (1) RAG applications are built as modular pipelines as shown in Figure 1(a), where a single query undergoes multiple rounds of LLM calls and retrievals, each of which serves different functions to improve overall output quality. For example, query transformation rounds (Gao et al., 2023a; Zheng et al., 2023; Peng et al., 2023; Asai et al., 2024a) generally occur before retrieval to refine the user’s query with LLMs (pre-retrieval generation). (2) RAG’s datastores are typically large, supported by recent works demonstrating that larger datastore consistently improves the performance of RAG applications (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]).
These characteristics create several significant challenges for efficient RAG inference, especially in latency-sensitive applications such as customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024), financial analysis (Malec, 2024; MyScale, 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024). First, the latency of RAG systems is inherently increased due to multiple rounds of LLM generation and retrieval. Second, although GPUs can speed up both LLM generation and retrieval, the combination of these processes poses significant memory requirements, often exceeding available GPU resources (see §3.1). Hence, it is expensive or even infeasible to fully host both operations on GPUs.
Consequently, in local or custom deployments, which are common for RAG applications with private or sensitive data, the retrieval datastore is often offloaded to CPU memory to alleviate GPU memory constraints. However, while CPU offloading resolves memory limitations, it significantly increases retrieval latency (see §3.1), diminishing overall system efficiency. This memory pressure is not limited to local deployments; even for serving scenarios in data centers, where many concurrent requests are processed together and GPUs are plentiful, managing large index sizes still presents a significant challenge for achieving high throughput with latency service level objectives (SLOs). Allocating additional GPU memory to the retrieval data store directly reduces the memory available for the KV cache of the LLM serving system, limiting the effective batch sizes (Kwon et al., 2023). In contrast, CPU retrieval would increase the per-request latency by a lot, highlighting the need for techniques that reduce GPU memory consumption while preserving low latency.
To address the latency challenge, several recent works have been proposed to accelerate RAG inference. These approaches include hardware accelerators for retrieval (Jiang et al., 2023d; Quinn et al., 2025), reuse of KV-cache (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024), and speculative document retrieval (Wang et al., 2024b; Zhang et al., 2024b). However, these works do not address the substantial memory demands associated with retrieval. Therefore, designing a memory-efficient system for low-latency RAG deployments is crucial.
In this paper, we argue that a promising direction to reduce GPU memory usage while maintaining low latency lies in improving the design of the retrieval index. Specifically, the inverted file index (IVF) (Sivic and Zisserman, 2003) reduces retrieval latency by pre-clustering data and limiting runtime search to relevant clusters. It offers a way to leverage GPU-accelerated search while not increasing memory consumption by dynamically transferring only the necessary clusters from the CPU to the GPU. However, this data transfer overhead is a key bottleneck, and needs to be minimized (Khan et al., 2024) (see §3.2).
Our proposal. To tackle latency and memory bottlenecks in RAG inference, we introduce a novel mechanism called lookahead retrieval, which predicts the subsets of IVF clusters that will likely be accessed during runtime and proactively transfers them from the CPU to the GPU before the retrieval stage.
The key observation behind lookahead retrieval is that there is substantial semantic overlap between queries before and after the pre-retrieval stage. Namely, IVF clusters used by the initial input query, which is available well before the retrieval stage, are also likely to be relevant to the actual retrieval query (see §3.3).
@@
Using this insight, we propose TeleRAG, an efficient inference system designed to optimize RAG latency while minimizing GPU memory consumption. TeleRAG employs lookahead retrieval to preemptively load relevant IVF clusters onto the GPU, effectively hiding CPU-GPU data transfer overhead during concurrent LLM generation. As illustrated in Figure 1(b), this approach significantly reduces retrieval latency without exceeding GPU memory constraints, enabling efficient execution of RAG applications. To maintain retrieval accuracy while adopting lookahead retrieval, TeleRAG simultaneously searches for clusters that were not prefetched (i.e., mispredicted) from the CPU and merges the results. TeleRAG also supports batch and multi-GPU inference.
A key challenge in deploying lookahead retrieval is determining the optimal number of IVF clusters to prefetch: Prefetching too many clusters increases data-transfer overhead, whereas prefetching too few could result in higher retrieval latency due to increased CPU processing. To address this, we propose a profile-guided approach coupled with an analytical model that dynamically determines the ideal prefetch amount based on pipeline characteristics and hardware configurations.
@@
  • Analyzing the correlation of the queries between the pre-retrieval generation and retrieval stages, revealing significant overlap in their corresponding IVF clusters.
@@
  • Proposing lookahead retrieval, which prefetches likely IVF clusters to the GPU, and hides CPU-GPU data transfer time during pre-retrieval generation.
@@
  • Developing TeleRAG, an efficient RAG inference system that integrates lookahead retrieval, resulting in significant acceleration of RAG with minimal GPU memory usage.
2. Background
Figure 2. Overview of RAG.
2.1. RAG
@@
RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text (Gao et al., 2023b; Asai et al., 2024c; Lewis et al., 2020). The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations (Mallen et al., 2023; Ram et al., 2023; Khandelwal et al., 2019) and generate contents based on up-to-date or private information (Izacard et al., 2023; Min et al., 2023).
RAG workflow. A basic RAG workflow involves several phases, including data store building, retrieval, and interactions with LLMs (Dixit, 2024; Ilin, 2023; Databricks, 2024; Wei, 2024; Vijeaswaran et al., 2024; Pathway, 2024; Gao, 2024). In order to build a data store, raw data in various formats is cleaned, converted to plain text, and divided into chunks. These chunks are then encoded into vectors using an embedding model and stored in a vector index, enabling efficient searching based on similarity. When a user provides a query, it is converted into a vector using the same embedding model, and the most similar chunks from the indexed database are retrieved. For a large-scale index, approximate algorithms such as the inverted file index (IVF) (Sivic and Zisserman, 2003) (see §2.2 for detail) are commonly used to accelerate the search process. The retrieved chunks, along with the original user query, are combined and given as a prompt to the LLM. The LLM then generates a response that relies on the information provided in the retrieved documents. This approach enables dynamic and informative responses in both single- and multi-turn dialogues (Lewis et al., 2020).
Modularity in modern RAG applications. However, naively integrating document retrieval into LLM generation can cause several issues. For example, the retrieval often struggles to find relevant content and may select irrelevant or mismatched chunks (Gao et al., 2023a; Zheng et al., 2023; Ma et al., 2023), and a single retrieval may not provide sufficient context, necessitating multiple retrieval rounds (Gao et al., 2023b).
To solve these issues, most state-of-the-art RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  (Gao et al., 2023b; Fan et al., 2024; Ilin, 2023; Gao, 2024; Databricks, 2024; Wei, 2024). Typically, they have the following types of steps (shown in Figure 2):
  • Pre-retrieval generation, used to assess whether retrieval is needed or to generate queries for retrieval. An example of a pre-retrieval technique is query transformation (Ma et al., 2023; Jagerman et al., 2023; Gao et al., 2023a; Zhou et al., 2022; Ye et al., 2023; Press et al., 2023; Peng et al., 2024; Zheng et al., 2023; Liu, 2022), which reformulates the original query to make it clearer and more suitable for the retrieval task.
  • Retrieval, used to identify relevant documents from the vector data store. This stage takes the output of pre-retrieval generation and generates data for the next stage.
  • Post-retrieval generation, generates response based on user query and retrieved documents. It can also perform additional process such as summarization (Jiang et al., 2023b; Kim et al., 2023b) or reranking (Zhuang et al., 2023; Stewart and Linsdell, 2024) on the retrieved documents.
  • Judgment, dynamically determines the execution flow. For example, it decides if more iteration is needed to enhance the response. Heuristics or LLMs can be used for judgement stage.
By proceeding through these functions, RAG applications can deliver more precise and contextually appropriate responses, significantly enhancing the capabilities of LLMs for various applications (Jin et al., 2024b; Gao et al., 2023b).
2.2. Vector Index and Inverted File (IVF)
$\text{k-argmin}_{i=0:N}d(x,y_{i}),$
where $D$ is the dimensionality of the vector determined by the embedding model, and $d$ denotes the distance function, which is typically the L2-norm or inner product (Johnson et al., 2019).
To improve search efficiency, the inverted file index (IVF) algorithm is widely used for large-scale vector databases due to its simplicity and effectiveness. IVF partitions the data store into clusters and restricts searches to the most relevant clusters, reducing the computational cost. To obtain the cluster assignment of each stored vector, it performs clustering algorithm such as $k$-means (Norouzi and Fleet, 2013) and partitions the database into $N_{c}$ clusters:
$\{C_{1},C_{2},\ldots,C_{N_{c}}\}=k\text{-means}(Y,N_{c}),$
$\text{Cluster}(y_{i})=\text{argmin}_{j=1:N_{c}}d(y_{i},c_{j}).$
$\{c_{j_{1}},c_{j_{2}},\ldots,c_{j_{L}}\}=L\text{-argmin}_{j=1:C}\ d(x,c_{j}).$
$k\text{-argmin}_{y_{i}\in\cup_{l=1}^{L}C_{j_{l}}}d(x,y_{i}).$
which involves sorting. In this way, IVF reduces search space and accelerates the retrieval process. Here, the hyperparameter $L$ from the first step is also referred as nprobe (Douze et al., 2024). When nprobe is larger, the IVF retrieval will be more accurate as it search more clusters. However, the retrieve latency is longer as more computation and data are needed.
Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs. Open-source libraries offer efficient GPU implementations (Johnson et al., 2019; Rapidsai, 2022). However, IVF does not reduce the memory requirement for the index since the data for all clusters must be stored. As a result, IVF still requires a substantial GPU memory footprint, which becomes a bottleneck when GPU memory capacity is the constraint of RAG systems. Moreover, recent work reported that increasing the size of the data store positively affects RAG application performance (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]), exacerbating this issue.
2.3. Use Cases and Motivation
In this paper, we aim to achieve low-latency RAG inference without driving up GPU memory consumption, which is relevant for both local and data center settings. This goal is critical for efficiently serving a wide range of real-world applications where users expect near-instantaneous responses, thus imposing strict latency SLOs. Examples include customer chatbots (Chemmagate, 2024; Akkiraju et al., 2024; Tung, 2024; AWS, [n. d.]), financial analysis (Malec, 2024; MyScale, 2024), autonomous driving (Cai et al., 2024; Yuan et al., 2024), and emergency medical diagnosis (Klang et al., 2024; GM and Dhar, 2024).
On the one hand, there is significant demand for RAG applications that involve user-specific private data in local settings (Wang and Chau, 2024; Lam et al., 2023; Chung et al., 2023; Wutschitz et al., 2023; Ghodratnama and Zakershahrak, 2023). In such scenarios, the available GPU resources are usually limited, and the system serves a single request at a time. Therefore, CPU offloading is necessary when using a large vector index, which introduces substantial latency as we will show in the next section.
On the other hand, for large-scale RAG services in data centers (per, [n. d.]; gen, [n. d.]; Asai et al., 2024a), the goal is to serve many concurrent user requests cost-effectively, necessitating high system throughput. However, maximizing LLM serving throughput is often bottlenecked by the GPU memory available for the KV cache, which dictates the maximum possible batch size (Kwon et al., 2023). If a large retrieval datastore is loaded onto GPUs, it consumes memory that could otherwise be used for a larger KV cache, thereby limiting throughput. Hence, falling back to slow CPU retrieval to save GPU memory is a common option (Jiang et al., 2025), but this is not desirable for latency-sensitive applications due to latency overhead.
In both cases, a straightforward way to reduce inference latency is scaling up GPU resources to allow both the LLM and datastore to reside in GPU memory, which is usually cost-prohibitive. Therefore, in this paper, we aim to design an efficient RAG inference system that can satisfy latency SLOs without requiring a significant amount of GPU memory.
3. Analyzing Latency of RAG Pipelines
@@
In this section, we analyze state-of-the-art RAG applications and identify their underlying systems challenges in achieving low latency. To conduct the analysis, we construct a 61 GB vector index with the Faiss library (Douze et al., 2024), and set the IVF clusters to 4096 following common practice (Asai et al., 2023). We use Llama-3-8B (AI@Meta, 2024) as the LLM, and run our analysis on an RTX 4090 (24 GB memory) and H100 (80 GB memory). §5.2 provides more details about our experimental setups.
3.1. End-to-end Latency of RAG Pipelines
@@
Figure 3. Latency breakdown of six RAG pipelines on NQ dataset (Kwiatkowski et al., 2019) with one H100 GPU. nprobe is 256.
Figure 4. The breakdown of memory consumption at GPU and CPU for two different strategies: CPU offloading and GPU-based retrieval. The dotted line indicates the memory capacity of a RTX4090 GPU, which is a common used GPU for local deployment.
We first analyze the end-to-end latency of six representative RAG pipelines (see §5.1 for details) in two scenarios: (1) the LLM runs on the GPU, while retrieval is offloaded to the CPU, minimizing GPU memory usage, and (2) both the LLM and the vector index reside in GPU memory, enabling GPU-based retrieval for lower latency. If the GPU’s memory is insufficient, CPU offloading is necessary.
We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details), to measure retrieval latency. Figure 3 shows the breakdown of end-to-end latency into LLM inference part and retrieval part, based on 1024 randomly sampled queries from the NQ dataset (Kwiatkowski et al., 2019). We observe that, in the CPU-based retrieval baseline (first scenario), the retrieval phase dominates the latency, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively. In contrast, GPU-accelerated retrieval in the second scenario dramatically reduces this bottleneck, accounting for only 10.5% and 28.3% of the latency in comparable configurations. On average, GPU retrieval is $5.96\times$ (batch size 1) or $3.87\times$ (batch size 4) faster than CPU retrieval, reducing overall latency by approximately $1.5\times$ or $1.8\times$, respectively. Thus, accelerating retrieval on the GPU is crucial for improving end-to-end latency.
However, GPU acceleration comes with a significant memory cost. Figure 4 shows the memory requirements for both GPU and CPU. As Figure 4 shows, putting both LLM weights and the retrieval index on the GPU requires around 77 GB of memory, which exceeds the capacity of consumer GPUs like the RTX 4090 with 24 GB. Thus, GPU acceleration on retrieval is often not feasible when running on a lower-end GPU or indexing with a large datastore.
Even on GPUs capable of holding the entire index in memory (e.g., H100), the index’s memory footprint limits throughput in data center serving scenario. Batched inference maximizes throughput, but it is bottlenecked by the KV cache capacity for LLM serving (Kwon et al., 2023). For the case of the example above, storing the index on the GPU leaves minimal memory for the KV cache (e.g., ~3 GB, or ~20k tokens worth), restricting batch size, while offloading the index to the CPU significantly increases the available memory (e.g., to ¿500k tokens worth). This issue is exacerbated since RAG typically involves long contexts (Jin et al., 2024a; Lu et al., 2024; Yao et al., 2024) and there is a trend of growing model and index sizes (Min et al., 2023; Khandelwal et al., 2019; Borgeaud et al., 2022; Hardt and Sun, 2023; Shao et al., [n. d.]), increasing memory pressure and necessitating offloading. However, CPU offloading is suboptimal for latency-sensitive applications as discussed above.
@@
Figure 5. Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
3.2. Opportunities of GPU-accelerated Retrieval with Limited Memory
A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on-demand for each query, leveraging the IVF index structure that narrows the search space (§2.2). While this method enables faster searching on the GPU, data fetching becomes the bottleneck.
In Figure 5, we compare the latency of the on-demand fetching system against CPU retrieval on an RTX 4090 GPU. We show three different nprobe values that determine the amount of data fetched. Overall, fetch time dominates latency due to the limited PCIe bandwidth between the CPU and GPU (32 GB/s). Although the GPU search is substantially faster than the CPU counterpart, the fetch overhead results in a higher end-to-end latency, which is about $3\%$ slower on average across nprobe values. Thus, to realize any meaningful speedup with this approach, the data fetching latency must be effectively hidden.
To hide data fetch costs, CPU-to-GPU transfer must precede retrieval, which requires predicting which data will be accessed. Fortunately, as we show next, the modern RAG pipelines offer a valuable hint as a query at the previous step.

### v3

@@ Abstract
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
1 Introduction
@@
Retrieval-augmented generation (RAG) has emerged as a powerful technique to enhance large language models (LLMs) by integrating them with external databases gao2023retrieval_rag_survey; asai2024reliable; lewis2020retrieval. During inference, RAG retrieves relevant content from external data sources, usually indexed as vector datastores, to mitigate issues such as hallucinations mallen2022not; ram2023context; khandelwal2019generalization and incorporate up-to-date or private information izacard2023atlas; min2023silo.
To provide accurate responses, modern RAG applications employ multiple rounds of LLM calls and retrievals for a single query gao2023retrieval_rag_survey; fan2024_rag_survey; advanced-rag-overview; modular-rag-survey; databricks-blog; azure-blog. As shown in Figure 1, RAG pipelines typically consist of four stages: (1) pre-retrieval generation: refines the initial user query to improve retrieval effectiveness. (2) retrieval: finds the relevant documents from datastore; (3) post-retrieval generation: generates a response based on the user’s query and retrieved documents; (4) judgment: Evaluates the output to determine the execution flow.
RAG datastores are typically large, as datastore size is strongly correlated with model accuracy min2023silo; khandelwal2019generalization; borgeaud2022improving; hardt2023test; shaoscaling. However, a brute-force search on such a large datastore is prohibitively slow and inefficient, as most data is irrelevant to the query. This drives the adoption of solutions like the Inverted File Index (IVF) ivf_original, which partitions the data into clusters and restricts the search to only the most relevant ones.
Figure 1: Typical pipeline stages of a RAG application.
While this partitioning effectively reduces the computational burden of the search, it does not solve the memory capacity challenge. To accelerate the IVF retrieval process with the GPU, one straightforward approach is to keep the entire datastore in the GPU memory (case A in Figure 2). As the size of a datastore is large (tens to thousands GB) quinn2025_iks, such a setting is expensive and often commercially impractical for local and custom RAG deployments handling private or sensitive data. Even in large data centers, where user requests are often batched, this approach challenges service level objectives (SLOs) because allocating GPU memory to the datastore reduces the LLM’s key-value cache (KV cache), limiting effective batch sizes kwon2023efficient. An alternative approach is to load the relevant data clusters from CPU to GPU at runtime. However, because bandwidth between the CPU and GPU is typically limited, this approach suffers from high transfer latency.
As a result, modern RAG systems jiang2025rago; shen2025hermes process retrieval on CPUs as they typically offer large and relatively cheap memory. However, CPUs are not as efficient as GPUs for highly parallelizable operations such as vector similarity search, and therefore using CPUs for search (case B in Figure 2) increases RAG system latency. For latency-sensitive applications such as customer chatbots vonage_reducing_rag_latency; akkiraju2024facts; ontinue_latency_ion_iq_chatbot, financial analysis harchworks_rag_financial; myscale_rag_trading, and emergency medical diagnosis klang2024assessing; apollo_medlm_rag, high latency can lead to poor user experience and even critical failures. To make matters worse, any latency increase of RAG systems has compounding effects due to multiple rounds of LLM generation and retrieval.
Figure 2: Illustration of TeleRAG’s lookahead retrieval mechanism and comparison to different RAG systems. TeleRAG prefetches relevant retrieval data from CPU to GPU, overlaps data transfer with the pre-retrieval stage, and accelerates retrieval with GPU–CPU cooperation.
Our proposal. In this work, we observe that there is a substantial semantic overlap between a user’s initial query and the query refined by an LLM during the pre-retrieval stage (see Figure 1). Since both queries represent the same core information need, their relevant IVF clusters are also likely to overlap significantly (see §3.3 for detailed analysis).
@@
Using this insight, we propose TeleRAG4), an efficient inference system that optimizes RAG performance by leveraging GPU retrieval acceleration while minimizing GPU memory consumption. TeleRAG employs lookahead retrieval to proactively load relevant IVF clusters onto the GPU, hiding the CPU–GPU data transfer overhead during concurrent LLM generation. As illustrated in Figure 2, our approach aims to significantly reduce retrieval latency without exceeding GPU memory constraints. To guarantee retrieval accuracy, TeleRAG complements this prefetching with a hybrid search: any clusters missed by the lookahead are simultaneously searched on the CPU, and the results are merged. In addition to prefetching, TeleRAG also utilizes on-GPU caching to further reduce redundant data transfers for frequently accessed clusters.
TeleRAG also optimizes for batch and multi-GPU inference (§4.2). To mitigate increased data transfer in batched scenarios, TeleRAG uses a prefetching scheduler to group similar queries, maximizing the overlap of their prefetched clusters. In multi-GPU settings, TeleRAG employs a cache-aware query scheduler that routes queries to the appropriate GPU to maximize the utility of its cached clusters.
@@
  • Analysis of the correlation of queries between the pre-retrieval generation and retrieval stages, revealing significant overlap in their corresponding IVF clusters.
@@
  • Lookahead retrieval, which prefetches likely IVF clusters to the GPU, and hides CPU–GPU data transfer time during pre-retrieval generation.
@@
  • TeleRAG, an efficient RAG inference system that integrates lookahead retrieval and supports high-performance multi-GPU inference through prefetching and cache-aware scheduling, resulting in significant acceleration of RAG with minimal GPU memory usage.
2 Background
2.1 Retrieval-Augmented Generation (RAG)
@@
RAG is a technique that enhances the capabilities of LLMs by integrating them with information retrieval to generate more accurate and relevant text gao2023retrieval_rag_survey; asai2024reliable; lewis2020retrieval. The core idea behind RAG is to augment the LLM with relevant information retrieved from a large corpus of documents, improving the LLM’s ability to answer questions without hallucinations mallen2022not; ram2023context; khandelwal2019generalization and generate contents based on up-to-date or private information izacard2023atlas; min2023silo.
Figure 3: Overview of RAG.
Modern RAG pipelines. In order to deliver precise and contextually appropriate responses, modern RAG models adopt a modular approach that employs multiple rounds of LLM calls and retrievals for a single query  gao2023retrieval_rag_survey; fan2024_rag_survey; advanced-rag-overview; modular-rag-survey; databricks-blog; azure-blog. Typically, they have the following types of steps (shown in Figure 3): (1) Pre-retrieval generation decides whether retrieval is needed or polishes queries before retrieval; for example, query transformation ma2023query; jagerman2023query; gao2022precise_hyde; zhou2022least; ye2023enhancing; press2022measuring; peng2024large; zheng2023take; llamaindex reformulates the original query to make it clearer and more suitable for retrieval. (2) Retrieval then identifies relevant documents for the query, taking the output of pre-retrieval generation and producing the evidence for the next stage. (3) Post-retrieval generation uses the user’s query and the retrieved documents to produce the final response; it can also include processes like summarization jiang2023longllmlingua; kim2023sure or reranking zhuang2023open; cohere-rerank. (4) Judgment component dynamically determines the execution flow (e.g., deciding whether further iteration is needed to enhance the response) using heuristics or LLMs.
2.2 Datastore, Inverted File Index (IVF), and Retrieval
Datastore. The datastore is processed by cleaning, chunking, and indexing to enable efficient retrieval. The raw data, in various formats, is first cleaned, converted to plain text, and divided into chunks. The chunks are then converted into vector embeddings using an embedding model such as Contriever izacard2021unsupervised. The document chunks, along with vector embeddings, are stored in a vector database, which enables efficient retrieval by searching for chunks based on vector similarities between the query and the chunk embeddings.
IVF-based retrieval. Many RAG deployments use the Inverted File Index (IVF) algorithm ivf_original for efficient vector search. IVF partitions the vectors into clusters based on similarity, each represented by a centroid. The search is then restricted to only the few clusters most relevant to the query, avoiding searching the entire datastore.
With clustering, an IVF search is a two-step process: (1) Cluster probing: The query is compared only to the cluster centroids to quickly identify the few most relevant clusters. (2) Searching: The system then performs a detailed search only within that small set of selected clusters and returns the final top-$k$ best-matching items.
A parameter called nprobe douze2024faiss controls the number of clusters to check in the first step. A larger nprobe increases accuracy at the cost of higher latency, as more data must be searched. Appendix A provides a formal description of IVF.
Latency of retrieval. Since the search process is highly parallelizable among each cluster and each vector, this search algorithm can be highly accelerated by GPUs. Furthermore, open-source libraries offer efficient GPU implementations johnson2019billion; rapidsai. However, the vector database in modern RAG deployments can reach tens to thousands of GB quinn2025_iks; sella2024flash as a result of the positive correlation of datastore size with accuracy min2023silo; khandelwal2019generalization; borgeaud2022improving; hardt2023test; shaoscaling. The large vector databases increase the latency of RAG pipelines significantly in both local setups and data centers.
Local setups typically consist of personal workstations or laptops with a single consumer-grade GPU (e.g., RTX4090 has 24 GB of memory), where latency is critical. Data center environments, on the other hand, care about both latency and throughput since they handle numerous simultaneous user queries. In both scenarios, GPU memory capacity becomes a performance limitation when working with vector databases. Consumer GPUs in personal setups lack sufficient memory to accommodate large vector databases, whereas data-center GPUs must sacrifice KV cache space (reducing throughput) to accommodate them kwon2023efficient. Consequently, vector databases are commonly stored in the CPU memory instead. Unfortunately, this approach increases retrieval latency whether IVF search operations are performed on the CPU or data must be transferred between CPU and GPU (as explored in §3).
3 Analyzing RAG Latency
@@
In this section, we analyze the system challenges in state-of-the-art RAG applications. We constructed a 61 GB vector index with the Faiss library douze2024faiss, and set the IVF clusters to 4096 following common practice asai2023self. Our analysis is conducted on an RTX 4090 (24 GB) and H100 (80 GB) using Llama-3-8B llama3modelcard. Further setup details are in §5.2.
3.1 End-to-end Latency of RAG Pipelines
@@
Figure 4: Latency breakdown of six RAG pipelines on NQ dataset kwiatkowski2019natural with one H100 GPU.
We first analyze the end-to-end latency of six RAG pipelines (§5.1) by comparing two scenarios: (1) LLM on GPU with retrieval on CPU (low GPU memory), and (2) both LLM and the vector index on GPU (low latency).
We set the nprobe to 256, a commonly used setting under this index’s configuration (see §5.2 for details). As shown in Figure 4 (using randomly sampled 1024 NQ queries kwiatkowski2019natural), the CPU-based retrieval phase is a major bottleneck, consuming 41.1% and 60.5% of total latency for batch sizes 1 and 4, respectively. In contrast, GPU-accelerated retrieval accounts for only 10.5% and 28.3% of the latency, respectively. On average, GPU retrieval is $5.96\times$ (batch size 1) and $3.87\times$ (batch size 4) faster than its CPU counterpart, reducing overall end-to-end latency by $1.5\times$ and $1.8\times$. Thus, accelerating GPU-based retrieval is crucial for reducing end-to-end latency.
However, GPU acceleration has a significant memory cost. For instance, in our setting, with a 61 GB index and a 16 GB Llama-3-8B model, GPU retrieval requires 77 GB of GPU memory, far exceeding the 24 GB of a common RTX 4090. This makes GPU-accelerated retrieval often unfeasible on lower-end GPUs or with large datastores.
Even on large GPUs (e.g., H100), storing the full index in memory limits serving throughput. High-throughput batched inference is bottlenecked by the LLM’s KV cache capacity kwon2023efficient. The index’s large memory footprint reduces the available memory for this KV cache, thereby limiting the effective batch size. This memory contention is exacerbated by RAG’s typically long contexts jin2024ragcache; lu2024turborag; yao2024cacheblend and the growing size of models and indexes min2023silo; khandelwal2019generalization; borgeaud2022improving; hardt2023test; quinn2025_iks, necessitating offloading.
@@
Figure 5: Latency breakdown of CPU-offload and runtime-fetch GPU retrieval, averaged over 512 random NQ queries.
3.2 GPU-accelerated Retrieval with Runtime Transfer
A straightforward approach to enable GPU retrieval with limited GPU memory is to fetch the necessary data from CPU to GPU on demand at runtime, leveraging the IVF index to narrow the search space. While this enables fast GPU search, the data transfer itself becomes the new bottleneck.
Figure 5 compares the latency of the runtime fetching system against CPU retrieval on an RTX 4090 GPU with three different nprobe values that determine the amount of data fetched. Overall, fetch time dominates latency due to the limited CPU–GPU PCIe bandwidth (32 GB/s). Although the GPU search is substantially faster, the fetch overhead results in a higher end-to-end latency ($\sim$3% slower on average). Thus, to achieve a meaningful speedup with this approach, the data-fetching latency must be effectively hidden.
To hide data-fetch costs, CPU-to-GPU transfers must occur before the retrieval stage, which requires predicting which data will be accessed. Fortunately, modern RAG pipelines offer a valuable hint: the query from the previous step.